# Fermion masses and mixings and charged lepton flavor violation in a 3-3-1 model with inverse seesaw

### v1

@@ Abstract
We propose a 3-3-1 model augmented with $A_{4}$ flavor symmetry to address fermion mass hierarchy, mixing patterns, and charged Lepton Flavor Violating (cLFV). The model incorporates an inverse seesaw mechanism to explain the small active neutrino masses through sterile neutrinos. Discrete symmetries, including $Z_{2},Z_{2}^{\prime},Z_{3},Z_{4},Z_{7}$ and $Z_{10}$ distinct mixing for the quarks and neutrinos . The model successfully predicts observed fermion masses and mixing patterns. Additionally, the constrained parameter space, particularly in the lepton sector, allows for potentially observable cLFV processes like $\mu\rightarrow e\gamma$.
@@ I Introduction
The $A_{4}$ discrete flavour group has been used widely in the literature, due to its high predictive power and since it provides a nice description of the SM fermion mass and mixing pattern [12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 39, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55]. We adopt the $A_{4}$ group as discrete flavor symmetry of the $SU(3)_{C}\times SU(3)_{L}\times U(1)_{X}$ 3-3-1 model, leveraging its unique three-dimensional irreducible representation to naturally accommodate the three fermion families. The spontaneous breaking of symmetries generates the SM fermion mass and mixing patterns. The absence of exotic electric charges in fermion sector allows for low-scale linear or inverse seesaw mechanism, generating tiny neutrino masses and potentialy testable sterile neutrino at the $SU(3)_{L}\times U(1)_{X}$ breaking scale. Our model employs the inverse seesaw mechanism to generate light neutrino masses, thus making the model testable at the colliders.
The remainder of the paper is organized as follows. In Section II, we present the proposed model. Section III analyzes the implications of the model for quark masses and mixings. Lepton masses and mixings are discussed in Section IV. The scalar and gauge sectors of the model are detailed in Section V. Section VI derives the constraints on the model parameter space imposed by charged lepton flavor violating decays. We conclude in section VII.
@@ II The model
We propose an extension of the economical 3-3-1 model with three right handed Majorana neutrinos, where the $SU(3)_{C}\times SU(3)_{L}\times U(1)_{X}$ gauge symmetry is supplemented by the $U(1)_{L_{g}}$ global lepton number symmetry and the $A_{4}\times Z_{2}\times Z_{2}^{\prime}\times Z_{3}\times Z_{4}\times Z_{7}\times Z_{10}$ discrete group. In this model, the scalar sector consists of two $SU(3)_{L}$ scalar triplets and several gauge singlet scalars, introduced to generate viable and predictive SM fermion mass matrices that are consistent with the low energy SM fermion flavor data. The scalar and fermionic spectra, along with their transformations under the $SU(3)_{C}\times SU(3)_{L}\times U(1)_{X}\times U(1)_{L_{g}}\times A_{4}\times Z_{2}\times Z_{2}^{\prime}\times Z_{3}\times Z_{4}\times Z_{7}\times Z_{10}$ group, are presented in Tables 1 and 2, respectively. The dimensions of the $SU(3)_{C}$, $SU(3)_{L}$ and $A_{4}$ representations shown in Tables 1 and 2 are indicated by boldface numbers, while the different $U(1)_{X}$ and $Z_{N}$ charges are written in additive notation. Note that a field $\psi$ transforms under the $Z_{N}$ symmetry with a corresponding $q_{n}$ charge as: $\psi\rightarrow e^{\frac{2\pi iq_{n}}{N}}\psi$, $n=0,1,2,3\cdots N-1$. We choose the $A_{4}$ symmetry because it is the smallest non-abelian group possessing triplet and three singlet irreducible representations (as detailed in Appendix A), thus allowing us to naturally accommodate the three families of SM fermions. Additionally, the $A_{4}$, $Z_{3}$, $Z_{4}$ and $Z_{7}$ symmetries select the allowed entries of the SM down-type quark mass matrices and determine the hierarchical pattern among these entries, which is necessary to correctly reproduce the the observed SM fermion mass and mixing structure. Besides that, the $Z_{2}$ and $Z_{3}$ symmetries naturally explain the small values of the electron and down quark masses. The $Z_{10}$ symmetry separates the $A_{4}$ scalar triplets participating in the quark and neutrino Yukawa interactions from the ones appearing in the charged lepton Yukawa terms, thus allowing to treat these sectors independently. Finally, the $Z_{2}^{\prime}$ symmetry is crucial for obtaining a nearly diagonal sterile neutrino mass matrix with non degenerate sterile neutrinos. In our model, the masses for the SM charged fermions lighter than the top quark arise from non-renormalizable Yukawa interactions, after the spontaneous breaking of the discrete symmetries. Consequently, the spontaneous breaking of the $A_{4}\times Z_{2}\times Z_{2}^{\prime}\times Z_{3}\times Z_{4}\times Z_{7}\times Z_{10}$ discrete group generates the observed pattern of SM fermion masses and mixing angles. The light active neutrino masses are generated through an inverse seesaw mechanism, where the smallness of its $\mu$ parameter is naturally explained by having higher dimensional Yukawa interactions involving gauge singlet scalar fields charged under the model’s various discrete symmetries. Upon spontaneous breaking of these discrete symmetries, small lepton number violating Majorana mass terms are generated, leading to the $\mu$ parameter in the inverse seesaw mechanism. It is worth mentioning that the $U(1)_{L_{g}}$ global lepton number symmetry is spontaneously broken down to a residual discrete $Z_{2}^{(L_{g})}$ by the vacuum expectation values (VEVs) of the $U(1)_{L_{g}}$ charged gauge-singlet scalars $\xi_{i}$ ($i=1,2,3$), $\varphi$ having a nontrivial $U(1)_{L_{g}}$ charge, as indicated by Table 1. The residual discrete $Z_{2}^{(L_{g})}$ lepton number symmetry, under which the leptons are charged and the other particles are neutral, prevents the appearance of interactions involving an odd number of leptons, which is crucial to guarantee the stability of the proton. The corresponding massless Goldstone boson arising from the spontaneous breaking of the $U(1)_{L_{g}}$ global symmetry, i.e., the Majoron, is a $SU(3)_{L}$ scalar singlet, which does not cause any problems in our model. Additionally, the fermions in our model do not exhibit exotic electric charges. Consequently, the electric charge assumes the following form
@@
$v_{\rho}\sim v_{\xi}\sim v_{\varphi}\sim\lambda^{5}\Lambda\ll v_{\chi}\sim\lambda^{2}\Lambda<v_{\sigma_{i}}\sim v_{S_{k}}\sim v_{\eta}\sim v_{\zeta}\sim v_{\phi}\sim v_{\Phi}\sim v_{\Xi}\sim\lambda\Lambda,\hskip 5.69046pti=1,2,3,\hskip 5.69046ptk=1,2,$ (13)
@@ III Quark masses and mixings
where $c_{1}$, $b_{n}$ ($n=1,2$),$\ a_{i}$,$\ g_{i}$ ($i=1,2,3$) are $\mathcal{O}(1)$ dimensionless parameters, assumed to be real, excepting $a_{1}$, which is taken to be complex. Notice that we take $a_{1}$ to be complex since the CP violating phase in the quark sector is associated with the quark mixing angle in the $1$-$3$ plane, as indicated by the Standard parametrization of the CKM quark mixing matrix. Besides that, notice that parameters and $v_{\rho}=246$ GeV the scale of electroweak symmetry breaking. The SM quark mass matrices of Eq. (24), imply that the quark mixing angles are generated from the up type quark sector.
@@
The experimental values of the quark masses [120, 121], mixing angles and Jarlskog invariant [122] are well reproduced, as indicated in Table 3, for the following values of the quark sector parameters:
$c_{1}$ $\simeq$ $-2.08887,\hskip 28.45274ptb_{1}\simeq 2.83883,\hskip 28.45274ptb_{2}\simeq 2.76616,\hskip 28.45274pta_{1}\simeq 3.62856,\hskip 28.45274pta_{2}\simeq 0.805245$
[…]
@@
As shown in Table 3, our model successfully fit for the ten physical observables in the quark sector. The symmetries inherent in our model enable us to successfully explain the SM quark mass spectrum and mixing parameters, with quark sector effective free parameters of order unity. The correlation between the quark mixing angle $\sin\theta_{13}$ and the Jarlskog invariant is depicted in Figure 1. In the left panel (a), we observe a heat-map for our model, revealing a strong correlation intensity in the $\sin\theta_{13}$-$J_{q}$ plane, with a correlation of $0.92$. This indicates that the Jarlskog invariant is highly sensitive to variation in $\theta_{13}$, aligning with the experimental observation of its small value. Conversely, variations in $\theta_{23}$ have a negligible impact on the invariant. Additionally, a significant correlation is observed in the $\sin\theta_{23}$-$J_{q}$ plane. By analyzing this heat map, we can refine our parameter space scan for our model of interest. As previously noted, the right panel (b) graphically illustrates the correlation between $\sin\theta_{13}$ and the Jarlskog invariant $J_{q}$. In this figure, the color represents the correlation in the $\sin\theta_{13}$-$\sin\theta_{23}$ plane.
@@ IV Lepton masses and mixings
$\tan\left(2\theta\right)\equiv t_{2\theta}=\frac{\sqrt{3}(b-1)\left((b+1)\kappa_{2}-2a\kappa_{1}\right)}{\kappa_{2}\left(-2a^{2}+b^{2}+1\right)+2\kappa_{1}((a+2)b+a)}$ (71)
@@
$a_{1}^{(\ell)}$ $\simeq 0.427,\quad a_{2}^{(\ell)}\simeq 1.026,\quad a_{3}^{(\ell)}\simeq 0.882$ (73)
$A$ $\simeq 31.65\pm 3.12\ \text{meV},\quad B\simeq 27.46\pm 2.24\ \text{meV},\quad C\simeq 20.79\pm 9.31\ \text{meV}$
$\theta$ $\simeq(-107.89\pm 3.33)^{\circ},\quad\phi\simeq(-29.83\pm 2.07)^{\circ}$
$\alpha$ $\simeq(-21.63\pm 1.15)^{\circ},\quad\gamma\simeq(-337.70\pm 8.85)^{\circ}$
@@
$\kappa_{1}=-0.655001,\quad\kappa_{2}=0.726924,\quad x=-0.695083+0.718929i$ (74)
As shown in Table 4, our model is consistent with the experimental data on lepton masses and mixing. It is worth mentioning that our model allows to successfully reproduce the experimental values of the six physical observables of the neutrino sector with only four effective parameters.
@@
$\theta^{(\ell)}_{12}(^{\circ})$ $33.4495\pm 2.16479$ $34.3\pm 1.0$ $33.44_{-0.74}^{+0.77}$ $31.4-37.4$ $31.27-35.86$
$\theta^{(\ell)}_{13}(^{\circ})$ $8.54137\pm 0.262994$ $8.58_{-0.15}^{+0.11}$ $8.57\pm 0.12$ $8.13-8.92$ $8.20-8.93$
@@
$\delta_{\text{CP}}(^{\circ})$ $-87.1246\pm 7.79677$ $216_{-25}^{+41}$ $197_{-24}^{+27}$ $128-359$ $120-369$
Table 4: Model and experimental values of the neutrino mass-squared diferences, leptonic mixing angles, and the $CP$violating phase. Experimental values are taken from Refs. [3, 123].

### v2

@@ Abstract
[…]
@@ I Introduction
The $A_{4}$ discrete flavour group has been used widely in the literature, due to its high predictive power and since it provides a nice description of the SM fermion mass and mixing pattern [12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 39, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 117, 118]. We adopt the $A_{4}$ group as discrete flavor symmetry of the $SU(3)_{C}\times SU(3)_{L}\times U(1)_{X}$ 3-3-1 model, leveraging its unique three-dimensional irreducible representation to naturally accommodate the three fermion families. The spontaneous breaking of symmetries generates the SM fermion mass and mixing patterns. The absence of exotic electric charges in fermion sector allows for low-scale linear or inverse seesaw mechanism, generating tiny neutrino masses and potentialy testable sterile neutrino at the $SU(3)_{L}\times U(1)_{X}$ breaking scale. Our model employs the inverse seesaw mechanism to generate light neutrino masses, thus making the model testable at the colliders.
The remainder of the paper is organized as follows. In Section II, we present the proposed model. Section III analyzes the implications of the model for quark masses and mixings. Lepton masses and mixings are discussed in Section IV. The scalar and gauge sectors of the model are detailed in Section V. Section VI contains an analysis and discussion of the flavor-changing neutral currents. Section VII derives the constraints on the model parameter space imposed by charged lepton flavor violating decays. We conclude in section VIII.
@@ II The model
We propose an extension of the economical 3-3-1 model with three right handed Majorana neutrinos, where the $SU(3)_{C}\times SU(3)_{L}\times U(1)_{X}$ gauge symmetry is supplemented by the $U(1)_{L_{g}}$ global lepton number symmetry and the $A_{4}\times Z_{2}\times Z_{2}^{\prime}\times Z_{3}\times Z_{4}\times Z_{7}\times Z_{10}$ discrete group. In this model, the scalar sector consists of two $SU(3)_{L}$ scalar triplets and several gauge singlet scalars, introduced to generate viable and predictive SM fermion mass matrices that are consistent with the low energy SM fermion flavor data. The scalar and fermionic spectra, along with their transformations under the $SU(3)_{C}\times SU(3)_{L}\times U(1)_{X}\times U(1)_{L_{g}}\times A_{4}\times Z_{2}\times Z_{2}^{\prime}\times Z_{3}\times Z_{4}\times Z_{7}\times Z_{10}$ group, are presented in Tables 1 and 2, respectively. The dimensions of the $SU(3)_{C}$, $SU(3)_{L}$ and $A_{4}$ representations shown in Tables 1 and 2 are indicated by boldface numbers, while the different $U(1)_{X}$ and $Z_{N}$ charges are written in additive notation. Note that a field $\psi$ transforms under the $Z_{N}$ symmetry with a corresponding $q_{n}$ charge as: $\psi\rightarrow e^{\frac{2\pi iq_{n}}{N}}\psi$, $n=0,1,2,3\cdots N-1$. We choose the $A_{4}$ symmetry because it is the smallest nonabelian group possessing triplet and three singlet irreducible representations (as detailed in Appendix A), thus allowing us to naturally accommodate the three families of SM fermions. Additionally, the $A_{4}$, $Z_{3}$, $Z_{4}$ and $Z_{7}$ symmetries select the allowed entries of the SM down-type quark mass matrices and determine the hierarchical pattern among these entries, which is necessary to correctly reproduce the observed SM fermion mass and mixing structure. In addition to that, the $Z_{2}$ and $Z_{3}$ symmetries naturally explain the small values of the electron and down quark masses. The $Z_{10}$ symmetry separates the $A_{4}$ scalar triplets participating in the quark and neutrino Yukawa interactions from the ones appearing in the charged lepton Yukawa terms, thus allowing one to treat these sectors independently. Finally, the $Z_{2}^{\prime}$ symmetry is crucial for obtaining a nearly diagonal sterile neutrino mass matrix with non-degenerate sterile neutrinos. In our model, the masses for the SM charged fermions lighter than the top quark arise from non-renormalizable Yukawa interactions, after the spontaneous breaking of the discrete symmetries. Consequently, the spontaneous breaking of the $A_{4}\times Z_{2}\times Z_{2}^{\prime}\times Z_{3}\times Z_{4}\times Z_{7}\times Z_{10}$ discrete group generates the observed pattern of SM fermion masses and mixing angles. The light active neutrino masses are generated through an inverse seesaw mechanism, where the smallness of its $\mu$ parameter is naturally explained by having higher dimensional Yukawa interactions involving gauge singlet scalar fields charged under the model’s various discrete symmetries. Upon spontaneous breaking of these discrete symmetries, small lepton number violating Majorana mass terms are generated, leading to the $\mu$ parameter in the inverse seesaw mechanism. It is worth mentioning that the $U(1)_{L_{g}}$ global lepton number symmetry is spontaneously broken down to a residual discrete $Z_{2}^{(L_{g})}$ by the vacuum expectation values (VEVs) of the $U(1)_{L_{g}}$ charged gauge-singlet scalars $\xi_{i}$ ($i=1,2,3$), $\varphi$ having a nontrivial $U(1)_{L_{g}}$ charge, as indicated by Table 1. The residual discrete $Z_{2}^{(L_{g})}$ lepton number symmetry, under which the leptons are charged and the other particles are neutral, prevents the appearance of interactions involving an odd number of leptons, which is crucial to guarantee the stability of the proton. The corresponding massless Goldstone boson arising from the spontaneous breaking of the $U(1)_{L_{g}}$ global symmetry, i.e., the Majoron, is a $SU(3)_{L}$ scalar singlet, which does not cause any problems in our model. Additionally, the fermions in our model do not exhibit exotic electric charges. Consequently, the electric charge assumes the following form
@@
[…]
In our model, the Yukawa terms contain several non-renormalizable interactions, as observed in Eqs. (6) and (7). These operators arise at high energies, where heavy mediators induce effective terms at low energies. They allow for an explanation of the fermion mass hierarchy without fine-tuning. Additionally, these operators may play a key role in generating light neutrino masses within the inverse seesaw mechanism. These operators have a general structure of the form
$\overline{\psi_{L}}H_{1}H_{2}\Psi_{R}\frac{H_{3}^{n}}{\Lambda^{n+1}},$ (8)
where $\psi_{L}$ and $\Psi_{L}$ are fermions, $H_{i}$ are scalar fields that acquire VEVs, generating the observed hierarchy in the fermion masses. The non-renormalizable operators could arise from a more fundamental theory with new scalar fields and heavy fermions. In particular, the effective operators in (6) and (7) may originate from the following renormalizable interactions
$\overline{\psi_{L}}H_{4}\tilde{\Psi}_{R},$ (9)
where the fermions $\tilde{\Psi}$ is heavy intermediate states that, once integrated out of the low-energy spectrum, induce the effective operators appearing in our model. The Yukawa operators of non-normalizable up and down quarks are has a possible ultraviolet origin of the non-renormalizable terms is that they are generated at low energies through the Feynman diagrams presented in Figure 1, where $H^{\prime}_{i}$ are mediator fields with masses on the order of the model cutoff scale $\Lambda_{\text{int}}$. Meanwhile, the non-renormalizable lepton Yukawa operators can be generated at low energies from the Feynman diagrams shown in Figure 2.
@@
$v_{\rho}\sim v_{\xi}\sim v_{\varphi}\sim\lambda^{5}\Lambda\ll v_{\chi}\sim\lambda^{2}\Lambda<v_{\sigma_{i}}\sim v_{S_{k}}\sim v_{\eta}\sim v_{\zeta}\sim v_{\vartheta}\sim v_{\phi}\sim v_{\Phi}\sim v_{\Xi}\sim\lambda\Lambda,\hskip 5.69046pti=1,2,3,\hskip 5.69046ptk=1,2,$ (15)
@@
Figure 1: Feynman diagrams that induce the non-renormalizable operators in quark sector.
Figure 2: Feynman diagrams that induce the non-renormalizable operators in lepton sector.
@@ III Quark masses and mixings
where $c_{1}$, $b_{n}$ ($n=1,2$),$\ a_{i}$,$\ g_{i}$ ($i=1,2,3$) are $\mathcal{O}(1)$ dimensionless parameters, assumed to be real, excepting $a_{1}$, which is taken to be complex. Notice that we take $a_{1}$ to be complex since the CP violating phase in the quark sector is associated with the quark mixing angle in the $1$-$3$ plane, as indicated by the Standard parametrization of the CKM quark mixing matrix. In addition, notice that parameter $v_{\rho}=246$ GeV corresponds to the scale of electroweak symmetry breaking. The SM quark mass matrices of Eq. (26), imply that the quark mixing angles are generated from the up type quark sector.
@@
The experimental values of the quark masses [122, 123], mixing angles and Jarlskog invariant [124] are well reproduced, as indicated in Table 3. Consequently, the fit for the quark sector was performed by minimizing a $\chi^{2}_{q}$ function, defined as,
$\chi^{2}_{q}=\sum_{i}\left(\frac{O_{i}^{\text{model}}-O_{i}^{\exp}}{\sigma_{i}^{\exp}}\right)^{2}$ (28)
where $O_{i}=\{m_{i},\sin\theta_{ij}^{(q)},J_{q}\}$, where $m_{i}$ are the masses of the quarks $(i=u,c,t,d,s,b)$, $\sin\theta_{ij}^{(q)}$ is the sine function of the quark mixing angles (with $j,k=1,2,3$ ) and $J_{q}$ is the quark Jarlskog invariant. The supra indices represent the experimental (exp) and theoretical (model) values, and the $\sigma$ are the experimental errors. Obtaining the following values for our best-fit point for the following values of the quark sector parameters:
$\begin{array}[]{l l l l l l l l}c_{1}\simeq-2.09\pm 0.13,&b_{1}\simeq 2.84\pm 0.14,&b_{2}\simeq 2.77\pm 0.13,&\left|a_{1}\right|\simeq 3.63\pm 0.18,&a_{2}\simeq 0.80\pm 0.03,\\ a_{3}\simeq 0.99\pm 0.03,&g_{1}\simeq-0.89\pm 0.04,&g_{2}\simeq 0.95\pm 0.06,&g_{3}\simeq-2.12\pm 0.07,&\arg\left(a_{1}\right)\simeq 23.35\pm 1.57^{\circ}.\end{array}$ (29)
As shown in Table 3, our model successfully fit for the ten physical observables in the quark sector. The symmetries inherent in our model enable us to successfully explain the SM quark mass spectrum and mixing parameters, with quark sector effective free parameters of order unity. The correlation between the quark mixing angle $\sin\theta_{13}$ and the Jarlskog invariant is depicted in Figure 3. In the left panel (a), we observe a heat-map for our model, revealing a strong correlation intensity in the $\sin\theta_{13}$-$J_{q}$ plane, with a correlation of $0.92$. This indicates that the Jarlskog invariant is highly sensitive to variation in $\theta_{13}$, aligning with the experimental observation of its small value. Conversely, variations in $\theta_{12}$ have a negligible impact on the Jarlskog invariant. A significant correlation is also observed in the $\sin\theta_{23}$-$J_{q}$ plane. By analyzing this heat map, we can refine our parameter space scan for our model of interest. As previously noted, the right panel (b) graphically illustrates the correlation between $\sin\theta_{13}$ and the Jarlskog invariant $J_{q}$. In this figure, the color represents the correlation in the $\sin\theta_{13}$-$\sin\theta_{23}$ plane.
Finally, to close this section, we concisely discuss the collider signatures of the exotic quarks in our model. The exotic up type $J_{1}$ and down $J_{n}$ ($n=2,3$) quarks can be produced in pairs via the gluon fusion mechanism and Drell-Yan annihilation. Once produced, they can decay in several ways: $J_{1}\to H^{+}d_{i}$, $J_{1}\to W^{\prime+}d_{i}$, while for the $J_{n}$ quarks, the decay modes include $J_{n}\to H^{-}u_{i}$, $J_{n}\to W^{\prime-}u_{i}$ ($i=1,2,3$). Here, $H^{\pm}$ are physical electrically charged scalars associated with the third component of the $SU(3)_{L}$ $\rho$ scalar triplet. The electrically charged scalars $H^{-}$ ($H^{+}$) can further decay into the SM-charged lepton (antilepton) and neutrino. Consequently, the pair production of exotic quarks at a proton-proton collider can produce a final state composed of two jets, opposite sign dileptons, and missing energy. Thus, observing an excess of events involving the SM background in the opposite sign dileptons, two jets, and missing energy can be a possible signal of support of this model at the LHC. A detailed analysis of the collider signatures of exotic quarks in theories with extended $SU(3)_{C}\times SU(3)_{L}\times U(1)_{X}$ is provided in [125]. A comprehensive study of the exotic quark production and decay modes at colliders in our model goes beyond the scope of this work and is deferred for a future publication.
@@ IV Lepton masses and mixings
Sterile neutrinos can be produced in pairs at the Large Hadron Collider (LHC) through Drell-Yan annihilation, mediated by a heavy $Z^{\prime}$ gauge boson. These sterile neutrinos mix with the light active neutrinos, enabling their decay into Standard Model (SM) particles. As a result, the final decay products include an SM-charged lepton and a $W$ gauge boson. Therefore, an excess of events in the dilepton final states, surpassing the SM background, could serve as a potential signature of this model at the LHC. Investigations into inverse seesaw neutrino signatures at colliders, as well as the production of heavy neutrinos at the LHC, have been conducted in, for example, in [126, 127, 128, 129, 130, 131]. A thorough investigation of the collider implications of our model is beyond the scope of this work and will be explored in future research.
@@
$\tan\left(2\theta\right)\equiv t_{2\theta}=\frac{\sqrt{3}(b-1)\left((b+1)\kappa_{2}-2a\kappa_{1}\right)}{\kappa_{2}\left(-2a^{2}+b^{2}+1\right)+2\kappa_{1}((a+2)b+a)}.$ (74)
@@
$a_{1}^{(\ell)}$ $\simeq 0.427,\quad a_{2}^{(\ell)}\simeq 1.026,\quad a_{3}^{(\ell)}\simeq 0.882,$ (76)
$A$ $\simeq 31.65\pm 3.12\ \text{meV},\quad B\simeq 27.46\pm 2.24\ \text{meV},\quad C\simeq 20.79\pm 9.31\ \text{meV},$
$\theta$ $\simeq(-107.89\pm 3.33)^{\circ},\quad\phi\simeq(-29.83\pm 2.07)^{\circ},$
$\alpha$ $\simeq(-21.63\pm 1.15)^{\circ},\quad\gamma\simeq(-337.70\pm 8.85)^{\circ},$
@@
$\kappa_{1}=-0.655001,\quad\kappa_{2}=0.726924,\quad x=-0.695083+0.718929i.$ (77)
As shown in Table 4, our model is consistent with the experimental data on lepton masses and mixing. It is worth mentioning that our model allows to successfully reproduce the experimental values of the six physical observables of the neutrino sector with only four effective parameters. Analogously to the quark sector (see Eq. (28)), the fit of the leptonic sector is obtained by minimizing the function $\chi^{2}_{\ell}$, which reproduces the values in Table 4 and the best-fit point from Eqs. (76) and (77).
@@
$\theta^{(\ell)}_{12}(^{\circ})$ $33.4495\pm 2.1648$ $34.3\pm 1.0$ $33.44_{-0.74}^{+0.77}$ $31.4-37.4$ $31.27-35.86$
$\theta^{(\ell)}_{13}(^{\circ})$ $8.54137\pm 0.26299$ $8.58_{-0.15}^{+0.11}$ $8.57\pm 0.12$ $8.13-8.92$ $8.20-8.93$
@@
$\delta_{\text{CP}}(^{\circ})$ $-87.1246\pm 7.7968$ $216_{-25}^{+41}$ $197_{-24}^{+27}$ $128-359$ $120-369$
Table 4: Model and experimental values of the neutrino mass-squared diferences, leptonic mixing angles, and the CP-violating phase. Experimental values are taken from Refs. [3, 132].
@@
Another observable predicted by our models is the effective Majorana neutrino mass parameter associated with neutrinoless double-beta decay $\left(0\nu\beta\beta\right)$. This process plays a pivotal role in investigating the Majorana nature of neutrinos, as it involves a nuclear decay that emits two electrons without any accompanying neutrinos. Its experimental observation would imply that neutrinos are Majorana fermions, violating the lepton number by two units and thereby pointing toward physics beyond the Standard Model.
The key parameter describing the amplitude of this process is the effective Majorana mass, $m_{ee}$, given by
$m_{ee}=\biggl{|}\sum_{i}U_{ei}^{2}\,m_{\nu_{i}}\biggr{|},$ (78)
where $U_{ei}$ are the elements of the PMNS matrix and $m_{\nu_{i}}$ are the masses of the light active neutrinos. The theoretical estimation of this value depends on the neutrino mass hierarchy.
In Figure 5 illustrates the correlation between the effective Majorana neutrino mass parameter $m_{ee}$ and the neutrmass squared difference $\Delta m_{21}^{2}$,the color scale representing the leptonic mixing parameter $\sin^{2}\theta_{23}^{(\ell)}$. The predictions in this work are derived under the assumption of a normal neutrino mass hierarchy and are consistent with the experimentally allowed ranges for neutrino oscillation parameters. The effective Majorana neutrino mass parameter $m_{ee}$ is predicted to lie in the range $0.1\,\mathrm{meV}\lesssim m_{ee}\lesssim 5.0\,\mathrm{meV}$, which is well below the current experimental upper bound of $m_{ee}\lesssim 50\,\mathrm{meV}$ set by KamLAND-Zen [133] based on the non-observation of the decay of $0\nu\beta\beta$ for ${}^{136}\mathrm{Xe}$, with a half-life constraint of $T_{1/2}^{0\nu}>2.0\times 10^{26}\,\mathrm{yr}$ at the confidence level 90%. Furthermore, our predictions are consistent with cosmological bounds from the Planck collaboration, which constrain the sum of neutrino masses to $\sum m_{\nu_{i}}<0.115\,\mathrm{eV}$ at 95% confidence level [4], taking into account our scenario of Eq. (76), we obtain a value in the following range $0.01\,\mathrm{eV}\lesssim\sum m_{\nu_{i}}\lesssim 0.06\,\mathrm{eV}$ The results presented here provide a robust framework to explore the parameter space of $m_{ee}$ within the context of normal hierarchy, highlighting the complementarity of neutrinoless double beta decay experiments and cosmological observations in probing the fundamental properties of neutrinos.
Figure 5: Correlation between the squared neutrino mass difference $\Delta m_{21}^{2}$ and the effective Majorana neutrino mass parameter $m_{ee}$. The color scale represents the variation of the leptonic mixing parameter $\sin^{2}\theta_{23}^{(\ell)}$, highlighting its influence on the distribution of $m_{ee}$.